We know the value of our test statistic and we have the formula to compute the p-value.

Test statistic: z∗=2.23

pvalue=P(z>z∗)

Now compute the p-value and enter the result rounded to exactly 4 decimal places.
Answer #1

We know that the p-value is defined as the probability to the right of respective statistic.

Now, given that our statistic z*=2.23.

To compute p-value, ie. P(Z>z*),

Where, Z follows standard normal distribution.

So, we have to find

P(Z>2.23)

=1-P(Z<=2.23)

=1-phi(2.23)

Where, phi is the distribution function of the standard normal variate.

=1-0.9871 (Value obtained from standard normal table)

=0.0129.

So, the answer is p-value=0.0129.